OpenStudy (anonymous):

how do you find the slope-intercept form of 7x=3y-12

OpenStudy (anonymous):

add 12 to both sides

OpenStudy (anonymous):

so 12+7x=3y

OpenStudy (anonymous):

then divide each term by 3

OpenStudy (anonymous):

so 4+7/3x=y

OpenStudy (anonymous):

and thats your answer

OpenStudy (chosenmatt):

Slope = 4.667/2.000 = 2.333 x-intercept = -12/7 = -1.71429 y-intercept = 12/3 = 4

OpenStudy (anonymous):

thanks that helped alot @stefandus @chosenmatt

OpenStudy (anonymous):

nppp

OpenStudy (chosenmatt):

if you need anything else just tag me!!! also can you give someone best response thanks!!

OpenStudy (anonymous):

is 4+7/3x=y in slope-intercept form @stefandus